# Trajectory-based models in strong-field physics

### v1

@@ 6.3 Nonsequential double ionization of molecules
[…]
@@
Semiclassical modelling has also been used to study molecular NSDI in Refs. [181, 182]. The study of Ref. [4] focuses on the NSDI process induced by a short ($6$ fs) laser pulse. Laser intensities corresponding to both tunneling and barrier suppression regimes (close to the threshold) were considered [181]. In the tunneling regime, two new mechanisms of double ionization have been revealed in Ref. [181]. The semiclassical simulations have revealed an anticorrelation pattern in distributions of electron momentum components along the polarization direction. This pattern is similar to the one experimentally observed for NSDI in Ar below the recollision threshold [165]. In contrast to this, relatively long laser pulses ($27$ fs) were employed in [182]. Again, laser intensities ranging from the tunneling to the barrier suppression regime were used in simulations. It is found that for intermediate intensities in the barrier-suppression regime, a characteristic square pattern prevails in the distributions of correlated electron momenta. The square pattern emerges due to the contribution of the delayed pathway of NSDI. This pathway corresponds to the situation where one of the electrons is ejected with a delay after the recollision. It is shown that for intermediate intensities the delayed pathway is dominated by the “soft" recollision.

### v2

@@ 6.3 Nonsequential double ionization of molecules
$\mathbf{r}_{10}\mathbf{F}\left(t\right)-V^{1}_{ne}-\int\frac{\left|\psi\left(\mathbf{r}\right)\right|^{2}}{\left|\mathbf{r}_{10}-\mathbf{r}^{~\prime}\right|}d\mathbf{r}^{~\prime}=I_{p1},$ (24)
@@
Semiclassical modelling has also been used to study molecular NSDI in Refs. [181, 182]. The study of Ref. [4] focuses on the NSDI process induced by a short ($6$ fs) laser pulse. Laser intensities corresponding to both tunneling and barrier suppression regimes (close to the threshold) were considered [181]. In the tunneling regime, two new mechanisms of double ionization have been revealed in Ref. [181]. The semiclassical simulations have revealed an anticorrelation pattern in distributions of electron momentum components along the polarization direction. This pattern is similar to the one experimentally observed for NSDI in Ar below the recollision threshold [165]. In contrast to this, relatively long laser pulses ($27$ fs) were employed in [182]. Again, laser intensities ranging from the tunneling to the barrier suppression regime were used in simulations. It is found that for intermediate intensities in the barrier-suppression regime, a characteristic square pattern prevails in the distributions of correlated electron momenta. The square pattern emerges due to the contribution of the delayed pathway of NSDI. This pathway corresponds to the situation where one of the electrons is ejected with a delay after the recollision. It is shown that for intermediate intensities the delayed pathway is dominated by the “soft” recollision.